# Know Unreported Roadway Incidents in Real-time: A Deep Learning Framework for Early Traffic Anomaly Detection

### v1

@@ 3 Data Preparation & Feature Engineering
We first discuss multiple data sources that are commonly available to use for early anomaly detection. Those data are most general in the sense they can be obtained to cover any geographic region and thus generally applicable to all locations. The data can be publicly available, or are offered by a number of data vendors with relatively mature technologies in the present market. In particular, we use incident reports, probe speed (segment level), and weather conditions for this study. Notably, all the information we use is gathered without the need to install additional detectors on the road network.
@@ 3.1 Incident Reports
Incident data can come from Waze and the State Departments of Transportation, which typically provide incident data feeds (real-time and historical). We select incident types that typically cause non-recurrent changes in traffic. Due to system discrepancies, the selected incident types vary slightly among various systems. The selected incidents mainly include accidents, hazardous weather, special events, and crashes. Those incident reports come with geographic locations and report time for each incident, which is translated to a binary indicator, $INC_{i}(t)$, denotes whether there is an incident report on link $i$ at time $t$.
@@ 3.2 Speed Data
Raw Traffic Speed: In our experiments, speed data comes from INRIX, which calculates speeds through real-time monitoring of probe vehicles. Similar data can also be obtained from other sources, e.g., HERE, AirSage, TomTom, or telematics data. We use four types of INRIX speed data: 1-minute granularity speed for private vehicles, trucks, and all vehicles, as well as 5-minute granularity speed for all vehicles. To describe the speed on a link, we primarily refer to the 5-minute granularity speed for all vehicles, as it exhibits a lower noise level and provides a comprehensive measurement across different vehicle types.
@@
where $\tau$ represents the set of upstream link segments within a certain distance range, and |.||.|| . | denotes the number of elements in the set. Equation (4) means that for a target link segment $i$, its slowdown speed is the difference between a target link and the speed of all its upstream links. A large slowdown speed indicates the presence of a sudden back-of-sequence slowdown in the spatial domain (Yao and Qian, 2020), and has a high correlation with the occurrence of incidents according to the incident shock-wave theory (Wirasinghe, 1978).
@@ 3.4 Data for experiments
As an experiment implemented in the later part of this paper, incident reports from two regions, Howard County, Maryland, and Cranberry Township, Pennsylvania, were collected for analysis. The data from Howard County spans 2022-2023, while data from Cranberry Township covers 2022-2024. Figure 4 shows the Traffic Message Channels (TMC) network (red lines) and the recorded incident locations (blue dots) for two regions.
@@ 4 Methodology
Our goal is to train an effective model in early anomaly detection/prediction. The model’s input consists of multi-source information over a past period, and the output is the anomaly status for a near future period. The incident reports are referenced for anomaly status labelling. To build an effective model, the following issues must be considered.
  • Rarity and randomness issues of the incident reports: Regarding rarity, it leads to a severe imbalance between anomaly and normality data samples, which would hinder models from learning non-recurrent patterns. On the other hand, randomness may result in inconsistent data distributions between the training and validation/testing sets, posing challenges to model generalization.
@@
We propose a framework combining several strategies to address the above issues. The overall picture of the framework is depicted in Figure 5, where each dashed boxes corresponds to a strategy we used. The framework first constructs a sub-graph for each link segment to establish the link-level anomaly prediction model. Then, by referencing and analyzing incident reports, we generate our anomaly labels (i.e., ground truth). With the inputs and labels in place, the data split module generates training, validation, and test data, ensuring no data leakage while maximizing the number of training samples. The model training section introduces the training techniques employed. In parallel with training, the model is calibrated to further enhance its performance. The details of each strategy will be introduced separately in this section.
Figure 5: High-level System Diagram
@@ 4.2 Label Denoising by Prior Knowledge
Using incident reports directly as anomaly labels can introduce significant noise and hinder the convergence process. This is due to the inherent issues within incident reports, such as delays, missing, and inaccuracies. Additionally, not all incidents have a tangible impact on traffic conditions. For example, an incident on a multi-lane highway during off-peak hours might not affect speed or flow, thus resulting in an undetectable incident. This may not be of interest to traffic operators if there is no impact to mobility or safety for the time being. Consequently, using incident reports directly as labels can cause substantial confusion between anomaly and normality labels, diminishing their significance and learnability. To ensure the labels are meaningful and learnable, existing research typically uses a manual process to filter reported incidents; however, this introduces subjectivity and limits their scalability. Most importantly, manually selecting incident reports does not address the issue of missing reports, potentially causing unreported anomalies to be labeled as normal. Therefore, we propose a prior knowledge-based method for labeling anomalies. Besides filtering out insignificant or false reports, unreported incidents are also evaluated and labeled.
Given that traffic flow data often requires additional detectors and is not widely available, we use the slowdown speed described in Equation (4) as the prior knowledge to infer whether an incident report corresponds to a traffic anomaly. Theoretically, as explained in Section 3.2, a significant traffic incident of our interest should cause a significant slowdown speed. Table 1 empirically demonstrates the strong correlation between significant slowdown speeds and the occurrence of incidents. For instance, in the first row of Table 1, data from the Howard County reveals that only 1.27% of timestamps on the I-70 westbound study segment have incident reports. However, of those timestamps associated with the top 3% of highest slowdown speeds, 9.99% have incident reports, and over 60% of the entire incident reports fall within these top 3% slowdown speed timestamps. Other study segments in Table 1 also exhibit similar characteristics, revealing a high correlation between significant slowdown speeds and the occurrence of incidents.
@@
Study Segment Region Total Incident Incident Reports in Overlap with Incident
Location Reports (%) Top Slowdown Speeds (%) Reports (%)
@@ 4.3 Ahead Labeling
The algorithm 1 mainly extracts common features from incident reports. However, as shown in Figure 2, incident reports often lack the early stage of incidents. While algorithm 1 can label some missing or delayed reports, they do not address the inherent delay issue in the reporting process. Therefore, we label anomalies starting from a few time steps prior to the reported incident/anomaly time to supplement the lack of early characteristics in the anomaly samples. For example, an anomaly from 7:30 a.m. to 8:30 a.m. can be ahead-labeled by 15 minutes to become from 7:15 a.m. to 8:30 a.m. The algorithm is shown in Algorithm 2. The setting of $\theta_{ahead}$ balances the inclusion of early features and false anomalies. The larger the $\theta_{ahead}$, the more early incident features are included, but it also increases the likelihood of normal samples being labeled as anomalies. Practically, this would need to be tuned for each road segment or to the traffic operator’s preference. In our experiment, we use $3$ as the default value of $\theta_{ahead}$, which is 15 minutes given the prediction interval is 15 minutes.
@@
Algorithm 2 Anomly Ahead Labeling
@@ 4.4 Multi-step Prediction & Data Splitting
Multi-step anomaly prediction/detection on past traffic conditions can be formulated as a sequence-to-sequence prediction task. To mitigate the issue of limited samples, we use a sliding window to increase sample numbers by partitioning the dataset into multiple subsections. As illustrated in Figure 6, suppose our study period is from 6 AM to 9 PM, and we use past 1 hour’s data to predict the anomaly status for the next half hour, with predictions made every five minutes. Then, the first sample of the day is using data collected between 6:00 AM-6:55 AM to predict the anomaly status between 7:00 AM-7:25 AM; the second sample is using data collected between 6:05 AM-7:00 AM to predict the anomaly status between 7:05 AM-7:30 AM, and so on. Using a sliding window can increase the number of training samples to mitigate the limited data issues.
@@ 4.5 Model Training
When $w_{ano}=1$, Eq.(9) is the binary cross-entropy (BCE), which is commonly used for binary classification training due to its differentiability. The purpose of weighting is to overcome the imbalance in datasets caused by the scarcity of anomaly samples. Without weighting, the predominance of normality samples would bias the model to consistently output zero. For the setting of $w_{ano}$, the default practice is to divide the number of samples whose label are 0 (normality samples) by the number of samples whose label are 1 (anomaly samples). Further adjustment is a trade-off between recall and precision (definitions refer to Figure 7). Increasing the weight helps to improve recall, while decreasing it helps improve precision.
@@ 4.6 Threshold Adaptation
The trained model outputs a number between 0 and 1, which can be interpreted as a probability of an anomaly; however, determining when to trigger an alert still requires a threshold. The practice of using 0.5 as a threshold in a balanced dataset assumes that the model can perfectly output values close to 0 and 1. However, due to the setting of $w_{ano}$ and the model’s inability to converge perfectly, the 0.5 threshold lacks theoretical basis and may perform poorly. Therefore, it is necessary to tune an appropriate threshold for best model selection and development purposes.
To prevent overfitting on the training data, threshold-tuning and model selection is performed on a validation set. Considering that incidents occur randomly, the distribution of anomaly samples in the validation set may differ from that in the training set. Therefore, we use the F1 score for model selection. The validation set is divided into two parts. The validation set 1 is used to find the threshold that maximizes the F1 score in validation set 1. With the tuned threshold, we can select the best model by optimizing the F1 score in validation set 2. Validation set 2 serves to simulate the test set. In this way, the model is selected based on a threshold that adapts to the model output, avoiding the issue of discarding well-performing models due to a fixed threshold.
@@ 5 Results and Discussion
This section is divided into two parts: the results of early anomaly detection (with possible prediction) and incident report comparison. In the early anomaly detection part, we use the generated anomaly label (ground truth) as a reference to assess our model’s ability to detect or predict significant anomalies in advance. In the incident report comparison part, we will compare our prediction results to raw incident reports (with inaccurate, insignificant, missing, or delayed reports).
@@ 5.1 Early Anomaly Detection/Prediction
Our experiments were conducted on the two road networks introduced in Section 3. We divided our training, testing, and validation sets in a 7:2:1 ratio in chronological order. We chose the periods that are relatively more challenging to predict: Predictions were made on weekdays, from 6:00 a.m. to 8:30 p.m., forecasting the anomaly status for the next half hour.
We first show the relationship between recall, precision, F1 score, and accuracy of the model as the threshold changes. Taking the Howard County I-70 segment as an example, Figure 8 shows the recall, precision, and F1 score for predicting anomaly status 5 minutes, 15 minutes, and 25 minutes ahead. The left side shows the results of the validation set, and the right side shows the results of the test set. From the figures, it can observed that the model’s performance on the validation set is almost identical to that on the test set. As the prediction period increases, the best F1 score gradually decreases. As the threshold increases, recall increases while precision decreases. Selecting an appropriate threshold allows us to achieve high accuracy and a high F1 score. The range for this threshold is approximately between 0.5 and 0.7. Figure 9 shows the results trained with 15-minute ahead labeling. Similarly, as the prediction period increases, the best F1 score decreases. As the threshold increases, recall improves while precision drops. Selecting a threshold between 0.4 and 0.6 can balance recall and precision, ensuring high accuracy. It is worth noting that the ground truth differs between the cases with and without ahead labeling, so directly comparing these evaluation metrics between Figure 8 and Figure 9 lacks significance.
@@
To validate the scalability of our method, we conducted experiments on 10 highway segments across two road networks. Tables 2 and 3 show the results. To avoid the interference of $\theta_{ahead}$ settings affecting the quality of the labels and thus reducing the persuasiveness of the results, Tables 2 and 3 use results without the ahead label. The effect of the ahead label will be discussed in the next chapter.
All the baseline models have been well-tuned and trained. The abbreviations represent the following:
@@
  • SVM, Support Vector Machine, trained using the Radial Basis Function (RBF) kernel. Due to the impracticality of the computation time required for upsampling, we used downsampling for training. We trained six SVM models for prediction models ranging from 5 minutes to 30 minutes, in 5-min time increments.
@@
  • GAN, Generative Adversarial Net (Goodfellow et al., 2014). We first used GAN to learn the characteristics of anomaly samples and generate more anomaly samples to overcome the sample rareness issue. SVM is then used for classification. Similar to SVM, we need to build six models.
@@
  • LightGBM, Light Gradient Boosting Machine (G. Ke et al., 2017). LightGBM is an efficient gradient boosting framework that uses tree-based learning algorithms. It is designed to be distributed and efficient, capable of handling large-scale data. For prediction models ranging from 5 minutes to 30 minutes, we trained six LightGBM models.
@@
  • Trans: Our model trained with Transformer (Vaswani et al., 2017).
  • GTrans: Our model trained with GraphTrans (Wu et al., 2021).
From the two tables, it is evident that the choice of model significantly impacts prediction performance. Overall, the results of Trans and GTrans used in our method generally rank first and second but show slight differences. GTrans performs slightly better on most segments, likely due to its ability to capture topological structures. GAN and LightGBM stand out among the other baseline models. Although both use SVM as the classifier, GAN’s overall performance is superior to SVM, demonstrating the significance of generating additional anomaly samples using GAN proposed by Lin et al. (2020). Compared to RF, LightGBM performs better, which aligns with LightGBM’s suitability for handling high-dimensional data. Notably, in the results for I-79 and I-76 in Cranberry, our model is outperformed by GAN in predicting anomalies at 20, 25, and 30 minutes. We believe this is because we use a shared decoder for all six predictions, whereas GAN trains six separate models. It can be observed that our model’s prediction performance at 5, 10, and 15 minutes is significantly better than GAN’s, likely because the shared decoder primarily learns to predict shorter time intervals rather than longer ones.

### v2

@@ 3 Data Preparation & Feature Engineering
We first discuss multiple data sources that are commonly available to use for early anomaly detection. Those data are most general in the sense they can be obtained to cover any geographic region and thus generally applicable to all locations. The data can be publicly available or offered by a number of data vendors with relatively mature technologies in the present market. In particular, we use incident reports, probe speed (segment level), and weather conditions for this study. Notably, all the information we use is gathered without the need to install additional detectors on the road network.
@@ 3.1 Incident Reports
Incident data can come from Waze and the State Departments of Transportation, which typically provide incident data feeds (real-time and historical). In our experiments, we use both Waze data and State DOT data, specifically from Pennsylvania DOT and Maryland DOT, depending on the region studied. We select incident types that typically cause non-recurrent changes in traffic. Due to system discrepancies, the selected incident types vary slightly among various systems. The selected incidents mainly include accidents, hazardous weather, special events, and crashes. Those incident reports come with geographic locations and report time for each incident, which is translated to a binary indicator, $INC^{i}(t)$, which denotes whether there is an incident report on link $i$ at time $t$.
@@ 3.2 Speed Data
Probe Vehicle Speed: In our experiments, speed data comes from INRIX, which calculates speeds through real-time monitoring of probe vehicles. Similar data can also be obtained from other sources, e.g., HERE, AirSage, TomTom, or telematics data. We use four types of INRIX speed data: 1-minute granularity speed for private vehicles, trucks, and all vehicles, as well as 5-minute granularity speed for all vehicles. To describe the speed on a link, we primarily refer to the 5-minute granularity speed for all vehicles, as it exhibits a lower noise level and provides a comprehensive measurement across different vehicle types.
@@
where $\tau$ represents the set of upstream link segments within a certain distance range, and |.||.|| . | denotes the number of elements in the set. Equation (4) means that for a target link segment $i$, its slowdown speed is the difference between a target link and the speed of all its upstream links. A large slowdown speed indicates the presence of a sudden back-of-sequence slowdown in the spatial domain (Yao and Qian, 2020) and has a high correlation with the occurrence of incidents according to the incident shock-wave theory (Wirasinghe, 1978).
@@ 3.4 Data for experiments
As an experiment implemented in the later part of this paper, incident reports from two regions, Howard County, Maryland, and Cranberry Township, Pennsylvania, were collected for analysis. The data from Howard County spans from 2022-02-14 to 2023-02-12, while data from Cranberry Township covers 2022-01-01 to 2024-01-31. Figure 4 shows the Traffic Message Channels (TMC) network (red lines) and the recorded incident locations (blue dots) for two regions. We select five segments from each region for our study (ten in total). The segment locations along with the number of incident reports for each segment are shown in Figure 5. It is important to note that the data from Cranberry Township spans two years, whereas the data from Howard County covers only one year, leading to a higher number of incidents reported in Cranberry Township.
@@
(a) Howard County, MD
(b) Cranberry Township, PA
Figure 5: Incident Report Count Per Segment
The INRIX speed data used in this study includes all-vehicle speeds at both 1-minute and 5-minute granularities, along with personal vehicle and truck speeds at a 1-minute granularity. All data with 1-minute granularity are imputed to a 5-minute granularity using Equation 3.
@@ 4 Methodology
Our goal is to train an effective model in early anomaly detection/prediction. The model’s input consists of multi-source information over a past period, and the output is the anomaly status for a near future period. The incident reports are referenced for anomaly status labeling. To build an effective model, the following issues must be considered.
  • Rarity and randomness issues of the incident reports: Regarding rarity, it leads to a severe imbalance between anomaly and normality data samples, which would hinder models from learning non-recurrent patterns. Regarding randomness, it may result in inconsistent data distributions between the training and validation/testing sets, posing challenges to model generalization.
@@
We propose a framework combining several strategies to address the above issues. The overall picture of the framework is depicted in Figure 6, where each gear pattern corresponds to a strategy we used. The framework first constructs a sub-graph for each link segment to establish the link-level anomaly detection/prediction model. Then, by referencing and analyzing incident reports, we generate our anomaly labels (i.e., ground truth) through prior knowledge and ahead-labeling. With the inputs and labels in place, the data split module generates training, validation (also including threshold-tuning set), and testing data, ensuring no data leakage while maximizing the number of training samples. The model training strategy includes the training techniques employed to deal with the anomaly rarity issue. In parallel with training, the alert threshold is calibrated to further enhance its performance. The details of each strategy will be introduced separately in this section.
Figure 6: Framework Overview
@@ 4.2 Label Denoising by Prior Knowledge
Using incident reports directly as anomaly labels can introduce significant noise and hinder the convergence process. This is due to the inherent issues within incident reports, such as delays, missing, and inaccuracies. Additionally, not all incidents have a tangible impact on traffic conditions. For example, an incident on a multi-lane highway during off-peak hours might not affect speed or flow, thus resulting in an undetectable incident. This may not be of interest to traffic operators if there is no impact on mobility or safety for the time being. Consequently, using incident reports directly as labels can cause substantial confusion between anomaly and normality labels, diminishing their significance and learnability. To ensure the labels are meaningful and learnable, existing research typically uses a manual process to filter reported incidents; however, this introduces subjectivity and limits their scalability. Most importantly, manually selecting incident reports does not address the issue of missing reports, potentially causing unreported anomalies to be labeled as normal. Therefore, we propose a prior knowledge-based method for labeling anomalies. Besides filtering out insignificant or false reports, significant anomalies, which are potential unreported incidents, are also evaluated and labeled.
Given that traffic flow data often requires additional detectors and is not widely available, we use the slowdown speed described in Equation (4) as the prior knowledge to infer whether an incident report corresponds to a traffic anomaly. Theoretically, as explained in Section 3.2, a significant traffic incident of our interest should cause a significant slowdown speed. Table 1 empirically demonstrates the strong correlation between significant slowdown speeds and the occurrence of incidents. For instance, in the first row of Table 1, data from Howard County reveals that only 1.27% of timestamps on the I-70 westbound study segment have incident reports. However, of those timestamps associated with the top 3% of highest slowdown speeds, 9.99% have incident reports, and over 60% of the entire incident reports fall within these top 3% slowdown speed timestamps. Other study segments in Table 1 also exhibit similar characteristics, revealing a high correlation between significant slowdown speeds and the occurrence of incidents.
@@
Study Segment Region Time Reported Incident Reports in Overlap with Incident
Location Incidents (%) Top Slowdown Speeds (%) Reports (%)
@@ 4.3 Ahead Labeling
The algorithm 1 mainly extracts common features from incident reports. However, as shown in Figure 2, incident reports often lack the early stage of incidents. While algorithm 1 can label some missing or delayed reports, they do not address the inherent delay issue in the reporting process. In addition, setting a fixed slowdown speed threshold may result in early-stage incidents—those that are developing but have not yet crossed the threshold—not being labeled. Therefore, we label anomalies starting from a few time steps prior to the reported incident/anomaly time to supplement the lack of early characteristics in the anomaly samples. For example, an anomaly from 7:30 a.m. to 8:30 a.m. can be ahead-labeled by 15 minutes to become from 7:15 a.m. to 8:30 a.m. The algorithm is shown in Algorithm 2. The setting of $\theta_{ahead}$ balances the inclusion of early features and false anomalies. The larger the $\theta_{ahead}$, the more early incident features are included, but it also increases the likelihood of normal samples being labeled as anomalies. Practically, this would need to be tuned for each road segment or to the traffic operator’s preference. In our experiment, we use $3$ as the default value of $\theta_{ahead}$, which is 15 minutes given the prediction interval is 15 minutes.
@@
Algorithm 2 Anomaly Ahead Labeling
@@ 4.4 Multi-step Prediction & Data Splitting
Multi-step anomaly prediction/detection on past traffic conditions can be formulated as a sequence-to-sequence prediction task. To mitigate the issue of limited samples, we use a sliding window to increase sample numbers by partitioning the dataset into multiple subsections. As illustrated in Figure 7, suppose our study period is from 6 AM to 9 PM, and we use past 1 hour’s data to predict the anomaly status for the next half hour, with predictions made every five minutes. Then, the first sample of the day uses data collected between 6:00 AM-6:55 AM to predict the anomaly status between 7:00 AM-7:25 AM; the second sample uses data collected between 6:05 AM-7:00 AM to predict the anomaly status between 7:05 AM-7:30 AM, and so on. Using a sliding window can increase the number of training samples to mitigate the limited data issues.
@@ 4.5 Model Training
When $w_{ano}=1$, Eq.(9) is the binary cross-entropy (BCE), which is commonly used for binary classification training due to its differentiability. The purpose of weighting is to overcome the imbalance in datasets caused by the scarcity of anomaly samples. Without weighting, the predominance of normality samples would bias the model to consistently output zero. For the setting of $w_{ano}$, the default practice is to divide the number of samples whose labels are 0 (normality samples) by the number of samples whose labels are 1 (anomaly samples). Further adjustment is a trade-off between recall and precision (definitions refer to Figure 8). Increasing the weight helps to improve recall, while decreasing it helps improve precision.
@@ 4.6 Threshold Adaptation
The model is optimized on the training set via a differentiable loss function, producing continuous output values ranging from 0 to 1. However, these continuous outputs themselves lack clear practical interpretation and can vary significantly across different segments, making them difficult to directly apply in real-world traffic management. To better meet practical needs, we introduce a threshold to determine when to trigger an alert. In other words, we need a threshold to convert the continuous outputs into meaningful binary classifications (0 or 1).
The practice of using 0.5 as a threshold might lead to suboptimal performance, particularly in imbalanced datasets. Additionally, traffic incidents exhibit seasonal and random characteristics—that is, their occurrence may be unevenly distributed over time. Therefore, we introduce a separate threshold tuning set. This dedicated dataset enables the optimization of threshold selection, ensuring that both the model and the chosen threshold generalize effectively across diverse incident distributions, thus improving model robustness. Our method for tuning thresholds and selecting models
Inputs: Training set $\mathcal{D}_{\text{train}}$, Threshold tuning set $\mathcal{D}_{\text{tune}}$, Number of epochs $E$
Outputs: Best model parameters $\theta^{*}$, optimal threshold $\tau^{*}$
initialize model parameters $\theta$
[…]
for $epoch=1$ to $E$ do
    Train the model $\theta$ on $\mathcal{D}_{\text{train}}$ using WBCE loss
[…]
    for $\tau=0.01,0.02,\ldots,0.99$ do
        Apply model $\theta$ to $\mathcal{D}_{\text{tune}}$ with threshold $\tau$
        Compute F1 score $\text{F}1_{\tau^{\prime}}$ under threshold $\tau$
[…]
           $\text{F}1_{\tau}\leftarrow\text{F}1_{\tau^{\prime}}$
           $\tau^{*}\leftarrow\tau$
        end if
    end for
    Apply model $\theta$ to $\mathcal{D}_{\text{val}}$ using threshold $\tau^{*}$
    Compute validation F1 score $\text{F}1_{val}$
    if $\text{F}1_{val}>\text{F}1_{best}$ then
        $\text{F}1_{best}\leftarrow\text{F}1_{val}$
        Save model parameters $\theta^{*}\leftarrow\theta$
    end if
end for
Algorithm 3 Threshold Adaptation
@@ 5 Results and Discussion
This section is divided into three parts: the results of label denoising, the results of early anomaly detection (with possible prediction), and incident report comparison. In the label denoising part, we compare our labeled anomalies with the incident report by analyzing speed deviation. In the early anomaly detection part, we use the generated anomaly label (ground truth) as a reference to assess different model’s ability to detect or predict significant anomalies. In the incident report comparison part, we compare our prediction results, as well as baseline prediction results, to raw incident reports (with inaccurate, insignificant, missing, or delayed reports).
5.1 Label Denoising Results
Given the extensive scale of the dataset, which spans over the years, expert-based manual labeling is infeasible. Therefore, we provide case studies to illustrate the effect of label denoising. As shown in Figure 9, on the training set of the I-70W segment, we demonstrate how our algorithm 1 labels additional anomalies in time periods where no incidents were reported. In figure 9, the black hatched areas represent the anomalies labeled by Algorithm 1, while the blue hatched areas indicate those labeled by Algorithm 2 (i.e., ahead-labeling). The blue line represents the traffic speed of the segment on the target day, and the yellow and green lines correspond to the recurrent speed under similar conditions (i.e., the nearby same day of the week and time of the day). The comparison shows that our algorithms can label abnormal traffic conditions that significantly deviate from the recurrent reference, even without incident reports. It is worth noting that although our algorithm is primarily based on spatial features—specifically, the detection of speed anomalies via the slowdown speed defined in Equation (4) —the labeled anomalies also exhibit significant deviation in the temporal dimension compared to the recurrent speed reference. The results show that the labeled anomalies are both temporal and spatial. Specifically, in cases (a), (b) and (d) of Figure 9, we did not assign labels during the final phase of the deviation, which may be due to the slowdown speed already falling below the threshold. However, it is acceptable since this study does not focus on predicting the duration of incidents. Meanwhile, this also helps us predict the end time earlier. Additionally, in case (c), although there appears to be no deviation along the temporal dimension, our algorithm still assigns labels, as the spatial impact of the anomaly has not yet dissipated.
In addition, the results highlighted in the blue hatched area further demonstrate the necessity of incorporating ahead-labeling. Since Algorithm 1 derives the abnormal slowdown speed threshold based on incident reports, it inevitably inherits a key limitation of such reports—the lack of early-stage abnormal samples. If we rely solely on the labels generated by Algorithm 1, many early abnormal patterns will be missing from the dataset. Consequently, models trained on this data may fail to effectively recognize the early stages of anomalies.
(a) 2022-02-18 Case
(b) 2022-02-21 Case
(c) 2022-02-23 Case
(d) 2022-07-01 Case
Figure 9: No Report, Anomaly Labeled
(a) 2022-03-04 Case
(b) 2022-05-18 Case
(c) 2022-05-31 Case
(d) 2022-06-06 Case
Figure 10: Reported, Anomaly Labeled
Similarly, Figure 10 presents the anomalies labeled by our algorithm for time periods with available incident reports. The red-shaded regions correspond to the durations of the reported incidents. It can be clearly observed that the anomalies labeled by our method are more temporally consistent with the deviation of speed.
Despite the absence of expert-labeled ground truth, which makes it difficult to directly quantify the overall performance of our label-denoising and ahead-labeling strategies, we observe that models trained with our generated labels consistently converge and achieve superior performance in real-time prediction tasks. In contrast, models trained on original incident reports often fail to converge. The learnability of the labels indirectly highlights the rationality of our labeling approach. A more detailed analysis is provided in Section 5.3.
@@ 5.2 Early Anomaly Detection/Prediction
In this section, we discuss the predictive performance of different models on our newly generated labels. Our experiments were conducted on the two road networks introduced in Section 3. We divided our training, testing, and validation sets in a 7:2:1 ratio in chronological order as explained in section 4.4. We chose the periods that are relatively more challenging to predict: Predictions were made on weekdays, from 6:00 a.m. to 8:30 p.m., forecasting the anomaly status for the next half hour. As discussed in Section 4.4, the problem can be formulated as a sequence-to-sequence (Seq2Seq) prediction task, for which the Transformer and its variants represent the current state-of-the-art. However, due to the inherent rarity and randomness of anomaly occurrences, effectively applying the Transformer model in our scenario necessitates further optimization and adaptation, as detailed in Sections 4.5 and 4.6.
We train the GraphTrans (Wu et al., 2021) using the introduced strategy and show the relationship between recall, precision, F1 score, and accuracy of the model as the threshold changes. Taking the Howard County I-70 segment as an example, Figure 11 shows the recall, precision, and F1 score for predicting anomaly status 5 minutes, 15 minutes, and 25 minutes ahead. The left side shows the results of the validation set, and the right side shows the results of the test set. From the figures, it can observed that the model’s performance on the validation set is almost identical to that on the test set. As the prediction period increases, the best F1 score gradually decreases. As the threshold increases, recall increases while precision decreases. Selecting an appropriate threshold allows us to achieve high accuracy and a high F1 score. The range for this threshold is approximately between 0.5 and 0.7. Figure 12 shows the results trained with 15-minute ahead labeling. Similarly, as the prediction period increases, the best F1 score decreases. As the threshold increases, recall improves while precision drops. Selecting a threshold between 0.4 and 0.6 can balance recall and precision, ensuring high accuracy. It is worth noting that the ground truth differs between the cases with and without ahead labeling, so directly comparing these evaluation metrics between Figure 11 and Figure 12 lacks significance.
@@
To further compare additional models, we conducted experiments on 10 highway segments across two road networks. Tables 2 and 3 show the results. To avoid the interference of $\theta_{ahead}$ settings affecting the quality of the labels and thus reducing the persuasiveness of the results, Tables 2 and 3 use results without the ahead label. The effect of the ahead label will be discussed in the next chapter. In addition to Transformer (Vaswani et al., 2017) and GraphTrans, we also evaluated models based on supervised learning algorithms that have demonstrated outstanding performance in AID tasks so far, including Random Forest, Light Gradient Boosting Machine, Support Vector Machine (SVM), and an SVM enhanced with data augmentation using a Generative Adversarial Network (GAN). All models were implemented with consideration for class imbalance and have been well-tuned and trained. The abbreviations represent the following:
@@
  • SVM, Support Vector Machine, trained using the Radial Basis Function (RBF) kernel, demonstrates superior performance in conventional AID (Yuan and Cheu, 2003). Due to the impracticality of the computation time required for upsampling, we used downsampling for training. We trained six SVM models for prediction models ranging from 5 minutes to 30 minutes, in 5-min time increments.
@@
  • GAN, Generative Adversarial Net (Goodfellow et al., 2014). Inspired by the idea proposed by Lin et al. (2020), We first use GAN to learn the characteristics of anomaly samples and generate more anomaly samples to overcome the sample rareness issue. SVM is then used for classification. Similar to SVM, we need to build six models.
@@
  • LightGBM, Light Gradient Boosting Machine (G. Ke et al., 2017). LightGBM is an efficient gradient boosting framework that uses tree-based learning algorithms. It is designed to be distributed, efficient, and capable of handling large-scale data compared to the XGBoost, which has shown superiority in anomaly detection (Parsa et al., 2020). For prediction models ranging from 5 minutes to 30 minutes, we trained six LightGBM models.
@@
  • Trans: Transformer (Vaswani et al., 2017) , trained with the optimization strategies described in Sections 4.5 and 4.6.
  • G-Trans: GraphTrans (Wu et al., 2021) , trained with the optimization strategies described in Sections 4.5 and 4.6.
From the two tables, it is evident that the choice of model significantly impacts prediction performance. Specifically, SVM demonstrates poor precision performance across all road segments, which means that it generates a substantial number of false positives. After incorporating GAN-based data augmentation, precision significantly improves despite a slight decrease in recall. It is noteworthy that, on our anomaly detection task, the observed decrease in recall following GAN augmentation differs from Lin et al. (2020)’s reported increase in detection rate. The observed performance gains from GAN-based augmentation in our anomaly detection task may be primarily attributed to the more balanced inclusion of both anomaly and normal samples during training. This balance potentially allows the model to capture the distribution of normal traffic conditions better. As a complementary analysis, training the SVM with a limited degree of upsampling—without introducing GAN-generated samples—could help clarify whether the improvements are mainly due to the increased representation of normal samples rather than the generative nature of GAN augmentation itself. Nevertheless, the positive impact of GAN augmentation on the overall model performance (F1 score) remains confirmed.
For tree-based models such as RF and LightGBM, LightGBM generally demonstrates a better balance between precision and recall while maintaining a high F1 score. In contrast, RF exhibits relatively poor recall performance, especially in regions such as I-695A, Howard County, and I-76W Cranberry Region. In these areas, the extremely low recall of RF leads to significantly worse overall performance compared to LightGBM and even SVM. This observation suggests that RF tends to make more conservative predictions, which aligns with the nature of bagging-based models—relying on majority voting, making the model more cautious in predicting positive cases. In contrast, LightGBM achieves superior overall performance due to its boosting-based architecture, which allows the model to iteratively correct misclassifications, especially for borderline samples, thereby improving its classification capability. By further comparing LightGBM with GAN, we found that LightGBM significantly outperformed GAN on the four road segments in Howard County. On other road segments, GAN performed slightly better, but the differences were not substantial.
Overall, Trans and G-Trans consistently outperformed the other models, primarily due to the superior data representation and learning capacity of their architectures. In particular, on the I-76W, I-76E, and US-19, Cranberry region, the Transformer model demonstrated a clear advantage in short-term prediction tasks. We attribute this performance to its effective temporal encoding mechanism. To further validate this hypothesis, we tested a variant of the Transformer model on the I-76E segment with the positional encoding removed. The results showed that the F1 scores dropped significantly to 0.41 and 0.29 for 5-minute and 10-minute predictions, respectively—substantially lower than those of the original model. This confirms that capturing and modeling temporal dependencies from the input data is essential. It also supports the necessity of our proposed strategy of incorporating temporal models into the prediction/detection framework. Compared with Trans, G-Trans performs slightly better on most segments, likely due to its ability to capture topological structures. Notably, in the results for I-79 and I-76 in Cranberry, Trans and G-Trans are outperformed by GAN in predicting anomalies at 20, 25, and 30 minutes. From the perspective of model architecture, this is probably because we use a shared decoder for all six predictions, whereas GAN trains six separate models. It can be observed that our model’s prediction performance at 5, 10, and 15 minutes is significantly better than GAN’s, likely because the shared decoder primarily learns to predict shorter time intervals rather than longer ones.